OpenStudy (anonymous):

During a research experiment, it was found that the number of bacteria in a culture grew at a rate proportional to its size. At 5:00 AM there were 4,000 bacteria present in the culture. At Noon, the number of bacteria grew to 4,600. How many bacteria will there be at midnight?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

want to do this the easy way?

OpenStudy (anonymous):

sure

OpenStudy (misty1212):

ok how many hours between 5:00 am and noon?

OpenStudy (anonymous):

7

OpenStudy (misty1212):

and it grew from 4,000 to 4,600 step one is to do \(\frac{4,600}{4,000}\) and reduce

OpenStudy (anonymous):

23/20

OpenStudy (misty1212):

ok so we can use \[P(t)=4,000\left(\frac{23}{20}\right)^{\frac{t}{7}}\]

OpenStudy (misty1212):

then figure out how many hours from 5:00 am to midnight, and use that for \(t\)

OpenStudy (anonymous):

So, 4000(23/20)^19/7? I got 5845.36

OpenStudy (misty1212):

hmm wait a sec

OpenStudy (misty1212):

ok looks good let me check the answer

OpenStudy (misty1212):

looks good to me \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]